Give me answer D,E


P1-9 Interest versus dividend income During the year just ended, Shering Distributors, Inc, had pretax earnings from operations of $490,000. In addition, during the year it received $20,000 in income from interest on bonds it held in Zig Manufacturing and received $20,000 in income from dividends on its 5% common stock holding dividend exclusion on its Tank Industries stock. a. Calculate the firms tax on its operating earnings only. b. Find the tax and the after-tax amount attributable to the interest income from c. Find the tax and the after-tax amount attributable to the dividend income from d. Compare, contrast, and discuss the after-tax amounts resulting from the interest e. What is the firms total tax liability for the year? Zig Manufacturing bonds. the Tank Industries, Inc., common stock income and dividend income calculated in parts b and c.
Answer #1

a.

Firm's tax on its operating earning

Operating Earning = $490,000

Tax on operating earning = 0.4 * 490000

Tax on Operating Earning = $196,000

b.

Interest Income from Zig Manufacturing Bonds = $20,000

Tax on Interest Income = 0.4 * 20000

Tax on Interest Income = $8,000

After tax amount of Interest Income = $12,000

c.

Dividend Income from Tank Industries = $20,000

Tax on Dividend Income after 70% dividend exclusion = 0.4 *(1 - 0.7) * 20000

Tax on Dividend Income = $2,400

After Tax amount of Dividend Income = $17,600

d.

Tax on Interest Income = $8,000

After tax amount of Interest Income = $12,000

Tax on Dividend Income = $2,400

After tax amount of Dividend Income = $17,600

Due to dividend exclusion clause of 70%, tax on dividend income is lower than tax on Interest Income, so even if before taxes income from interest and dividend are same, but tax on them makes a substantial difference in after tax values

e.

Firm's Total Tax Liability = 196000 + 8000 + 2400

Firm's Total Tax Liability = $206,400